the area of a rectangle is 6x^2-31x+35 square innches.
what is the length and width of the rectangle?

Found 2 solutions by mananth, Alan3354:
Answer by mananth(16949)

the area of a rectangle is 6x^2-31x+35 square innches.
what is the length and width of the rectangle?
..
6x^2-31x+35
6x^2-21x-10x+35=0
3x(2x-7)-5(2x-7)=0
(2x-7)(3x-5)=0
x= 7/2 OR x= 5/3
...
The dimensions are = 3.5 by 1.66 inches

Answer by Alan3354(69443)

Not sufficient info.
It could be 1 by 6x^2-31x+35, or 5 by (6x^2-31x+35)/5, or any of an infinite number of combinations.